Eliminate the parameter t.


x = 4 csc t - 4, y = 4cot t - 3


1
Expert's answer
2021-05-27T09:37:27-0400

x = 4 "\cosec t" - 4 ......................Equation(1)

y = 4 "\cot t" - 3 .......................Equation(2)



We need to eliminate t, so from Equation (1) and (2) we have


"\cosec t" = "\dfrac{x + 4}{4}" ........................Equation(3)


"\cot t" = "\dfrac{y+3}{4}" .........................Equation(4)


We know that


"\cosec^2 t - \cot^2 t = 1"


From Equation (3) and (4) we have


"(\dfrac{x + 4}{4})^2" - "(\dfrac{y+3}{4})^2" = 1


Hence, the parameter t has been eliminated.